Hi I need help with a couple of problems. I need to solve these system of equations by graphing and I don't know how. The problem is x+y=5
2y=x-2
that is the same problem. I just need to know how to show the steps to solve it and graph it.

Answer by jb007mi6(53)

so you have y=-x+5
y=x/2-2
so just graph the lines...and where they intersect is your solution
so since both of these equations =y you can set them equal to eachother like so...
-x+5=x/2-2
so now i would clear the fraction
-2x+10=x-4
add 2x and 4 to both sides
14=3x
divide both sides by 3 and get x=14/3
and now, you need to solve for y lets take the first equation since there's no fractions
y=14/3+15/3
so y = 29/3
so your solution is (14/3,29/3)